OpenStudy (anonymous):

A number exceeds its 3/5th by 16. What is the number?

OpenStudy (anonymous):

x-3/5x=16 x=40

OpenStudy (anonymous):

could you plz provide some explanation

OpenStudy (anonymous):

if the no is supposed to b x, then it is greater than its 3/5th by 16, mean the difference of the no n its 3/5th is 16

OpenStudy (amistre64):

i assume 3/5th is a percentage of some number that is 16 greater than the number?

OpenStudy (anonymous):

A number exceeds its 3/5th by 16

OpenStudy (amistre64):

n is greater than its (3/5)% by 16

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

though m not sure either :)

OpenStudy (amistre64):

I think its the same answer; but lets see :) n(1+3/5) = n+16

OpenStudy (amistre64):

8n --- = n+16 5 8n = 5n + 80 3n = 80 n = 80/3 = 2.666; maybe it is, maybe it aint lol

OpenStudy (anonymous):

counter check it now:)

OpenStudy (anonymous):

see if the 3/5% exceeds by 16?

OpenStudy (amistre64):

26.666 ; decimated in the wrong spot

OpenStudy (amistre64):

80 3 80 ---- * -- = --- = 16 3 5 5

OpenStudy (amistre64):

i dount i read the question right to begine with :)

OpenStudy (anonymous):

:)